OpenStudy (anonymous):

The second angle of a triangular building lot is three times as large as the first. The third angle is 10° more than the sum of the other two angles. Find the measure of the third angle. please show all work!

OpenStudy (anonymous):

the angles are x, 3x, (x+3x+10)=4x+10. The sum must be 180 x+3x+4x+10=180 8x=170 x=85/4 The third angle was 4x+10 4(85/4)+10=95

OpenStudy (anonymous):

Let the three angles be <1, <2, <3. We know that the sum of angles of a triangle should be 180. So eqn.1 is <1 + <2 + <3 = 180 -----(1) <2=3*<1----(2) <3=10+<1+<2 ------(3) Now solve the equations (1),(2),(3)